Difficult Addition


Arjun is trying to add two numbers. Since he has learned addition recently, an addition which requires a carry is difficult for him. Your task is to print Easy if the addition does not involve a carry, Otherwise print Hard.


Input

The first line of input contains two space separated integers A and B.


Output

If the addition does not involve a carry,print Easy, otherwise print Hard.


Explanation

When calculating 229 + 390, we have a carry from the tens digit to the hundreds digit, so the answer is Hard.


Sample Input1

229 390

Sample Output1

Hard


Sample Input2

123456789 9876543218

Sample Output2

Easy


1
Expert's answer
2022-02-25T07:53:03-0500
l=0
while(l!=2):
    s = (str(input("Enter two numbers separated by space: "))).split(" ")
    l = len(s)


Flag=0
for r in range(0,len(s[0])):
    
    t1 = int(s[0][r])
    t2 = int(s[1][r])
    if((t1+t2)>9): Flag=1
if(Flag):   print("HARD")
else:       print("EASY")